Be sure to answer all parts.

You are given solutions of HCl and NaOH and must determine their concentrations. You use 82.6 mL of NaOH to titrate 100. mL of HCl and 12.4 mL of NaOH to titrate 50.0 mL of 0.0782 M H2SO4. Find the unknown concentrations.
____ M NaOH
____ M HCl
Be sure to answer all parts.

You are given solutions of HCl and NaOH and must determine their concentrations. You use

82.6 mL

of NaOH to titrate 100. mL of HCl and

12.4 mL

of NaOH to titrate 50.0 mL of

0.0782 M H2SO4.

Find the unknown concentrations.
M NaOH
M HCl
Answer #1

Moles of H2SO4 used = Volume in Liters * Molarity = 0.050 L * 0.0782 M = 0.00391 mole

Moles of NaOH reacted with H2SO4 = 2 * 0.00391 mole = 0.00782 moles

So, Concentration of NaOH = Moles / Volume in Liters = 0.00782 mole / ( 0.0124 L ) = 0.63 (M)

For titration between HCl and NaOH:

Moles of HCl = Moles of NaOH used = Volume in Liters * Molarity of NaOH solution = 0.0826 L * 0.63 M = 0.052 mole

Concentration of HCl = 0.052 mole/ 0.100 L = 0.52 M